# Consumption Partial Insurance

### v3

@@ Abstract
We measure the extent of consumption insurance to income shocks accounting for high-order moments of the income distribution. We derive a nonlinear consumption function, in which the extent of insurance varies with the sign and magnitude of income shocks. Using PSID data, we estimate an asymmetric pass-through of bad versus good permanent shocks – 17% of a $3\sigma$ negative shock transmits to consumption compared to 9% of an equal-sized positive shock – and the pass-through increases as the shock worsens. We find similar patterns for transitory shocks among the least wealthy. Households are willing to sacrifice more than 1/8 of lifetime consumption to eliminate tail income risk. Our results are consistent with surveys of consumption responses to hypothetical events and suggest that tail risk matters substantially for consumption.
@@ 1 Introduction
In a seminal paper, Blundell et al. (2008) –henceforth, BPP– introduce an empirical methodology to measure the pass-through of income shocks of varying persistence into consumption. Using Panel Study of Income Dynamics (PSID) and Consumer Expenditure Survey (CEX) data over 1980–1992, BPP find evidence of partial insurance to permanent shocks and full insurance to transitory ones. BPP rely on a workhorse covariance between income and consumption growth, namely a second moment of their joint distribution. The idea is that the extent to which income growth (driven by income shocks) varies with consumption growth reflects the strength of the transmission of income shocks into consumption.
In a separate strand of literature, Guvenen et al. (2021) use data from the U.S. Social Security Administration and establish two new important facts about the distribution of unexplained income growth, namely the distribution of income shocks. First, income growth is very negatively skewed, i.e., it has a long left tail, so there are far more people in the data who experience large negative than large positive shocks. Second, income growth exhibits excess kurtosis, i.e., fat tails, so there are far more people in the data who face either small or large income shocks, than people who face moderate ones, relative to a Gaussian density. These are now benchmark facts about income dynamics, observed in many countries (e.g., De Nardi et al., 2021, compare the U.S. and the Netherlands).
If most shocks away from the mean are large and bad shocks, their effect on consumption might, intuitively, differ from that of small/moderate shocks. In other words, left skewness and excess kurtosis, which are statements about the third and fourth moments of the income distribution, respectively, may matter for the pass-through of shocks into consumption. This is precisely what we investigate in this paper. Specifically, we measure the degree of consumption insurance to income shocks of varying persistence, accounting for higher-order moments of their distribution (i.e., tail shocks). This is in contrast to the methodology in BPP, which revolves around the second moments of the joint income-consumption distribution. The second moment is a poor summary of the tails and, thus, abstracts from the modern higher-order facts that characterize income dynamics.
@@
Third, we generalize the method of measuring consumption insurance to explicitly permit a role for tail income risk. BPP log-linearize the policy rule of a lifecycle consumption/savings problem to obtain a consumption function that depends linearly on income shocks. Our first and second contributions above revolve firmly around such a linear framework. We argue (as also does Carroll, 2001) that linearity is a poor approximation to behavior when income is subject to tail shocks. We then derive a higher-order (quadratic) consumption function that allows for nonlinear transmission of shocks. We characterize and empirically measure the consumption pass-through of income shocks; in this environment, the degree of partial insurance depends on the magnitude and sign of shocks. While a nonlinear consumption function has been estimated before, most notably in Arellano et al. (2017), identification in our case is far less data demanding relative to them: we need second moments of consumption and up to fourth moments of income, rather than observing the entire consumption-income joint distribution. This has certain important advantages, which we discuss below.
Fourth, we estimate the welfare costs of tail income risk by calculating what fraction of lifetime consumption a household is willing to give up to avoid tail risk. De Nardi et al. (2019), Busch and Ludwig (2021), and Guvenen et al. (2024) provide similar calculations through structural models that fully specify the insurance channels at the nexus of income and consumption. By contrast, we remain agnostic about specific insurance mechanisms; we estimate an ‘aggregate’ of all partial insurance available to households, and then use a parametric utility function to estimate the welfare costs of tail risk.
@@
First, permanent and transitory shocks to disposable household income are negatively skewed, with skewness coefficients at $-0.95$ and $-1.31$, respectively, and highly leptokurtic, with kurtosis coefficients at $44.39$ and $49.23$, respectively. Left skewness implies that a negative shock is typically more unsettling, more hurtful than a positive shock, because it is further away from the zero mean. Excess kurtosis implies that more shocks are concentrated either in the middle of the distribution or far out in the tails, relative to the shoulders. Taken together, large left skewness and excess kurtosis suggest that the realizations of income shocks tend to be either close to their zero mean or far out in the left tail; that is, income shocks for most households are either small/average or quite bad.
@@
Second, in estimating BPP in the recent years, we find full insurance to transitory shocks and a small pass-through of permanent shocks at 0.15 versus 0.64 in BPP. A 10% permanent income cut thus reduces consumption by only 1.5%, vis-à-vis 6.4% in BPP. In other words, we find much larger partial insurance to permanent shocks in the recent years than originally found in 1980–1992.222Using the recent PSID data but a different method, Arellano et al. (2023) also find a low pass-through of permanent shocks. We discuss this point further when we present our empirical results. We attribute one third of the difference from BPP to the consumption imputation from the CEX that BPP use. The imputation brings in large amounts of measurement error, inflating the covariance between income and consumption growth. Another third of the difference arises because the modern PSID provides data biennially (versus annually previously), which tends to mute higher frequency shocks and their transmission into consumption. The last third reflects the different calendar times.
@@
Third, introducing the third and fourth moments of income and consumption growth into BPP’s linear consumption setting does not change our estimates of partial insurance. This is counterintuitive in light of the literature that highlights, in multiple settings, the large welfare implications of tail risk (e.g., Barro, 2009; De Nardi et al., 2019). One would thus expect that targeting higher-order moments would reduce the degree of partial insurance, which is not what we find. This result is suggestive of a potential misspecification in the linear consumption function, which motivates our generalization of the methodology through the introduction of a higher-order consumption specification subsequently.
@@
Fourth, in estimating our quadratic consumption function, we find an asymmetry in the pass-through of bad versus good shocks, and of small versus large ones. In this environment, the pass-through depends on the magnitude and sign of shocks; and targeting higher-order income moments enables the identification of the transmission parameters. We draw three main conclusions. First, bad permanent shocks are more hurtful than equal-sized good ones because they have larger transmission rates into consumption – a three standard deviations negative shock (a 50% permanent income cut) has a pass-through of $0.17$ (17% of the shock transmits into consumption), while a positive shock of similar magnitude has a pass-through of $0.09$. Second, the pass-through of bad shocks increases with the severity of the shock. Bad permanent shocks are thus double hurtful: not only are they more extreme than good ones (they lie further out in the left tail), but they also have increasingly larger transmission rates as their magnitude worsens. Third, transitory shocks provide mixed evidence. They are fully insured in the full sample, at least from a statistical point of view, although the point estimates of their transmission parameters suggest that large transitory windfalls shift consumption considerably, consistent with quasi-experimental evidence of large responses to transitory gains (e.g., Fagereng et al., 2021). Among low wealth households, however, transitory shocks shift consumption significantly, and negative shocks pass through at increasingly larger rates than positive ones. Overall, the results are consistent with an environment in which households save a good income shock but dissave in the presence of a bad one. As there is clearly a limit to how much one can dissave or borrow, a large negative shock ends up affecting consumption by relatively more.
@@
The asymmetric transmission of bad versus good permanent shocks intensifies with the age of the household. While older households insure average permanent shocks better than younger ones, older households have larger transmission rates of extreme shocks. Severely bad permanent shocks thus cause more harm among older than younger households. Guvenen et al. (2021) show that income skewness becomes more negative with the age of the earner – there is more room for income to fall as older workers have higher incomes; we find a similar pattern in our data too. In turn, we document that there is also more room for consumption to fall, so bad shocks are double hurtful among the old. Overall, the results from our quadratic specification are indicative of limited insurance to left tail income risk. Moreover, if the econometrician assumed a linear consumption function, she would overestimate the true extent of insurance to it, i.e., she would infer excessive insurance.
@@
The seminal paper in the first branch, namely BPP, sparked a large volume of research. [^3] Hall and Mishkin (1982), Deaton and Paxson (1994) and Blundell and Preston (1998) characterize the empirical distribution of consumption and income, so they are early precursors in this literature. Blundell et al. (2016) extend BPP to an environment with wage risk to measure the contribution of family labor supply to partial insurance. Theloudis (2021) extends this further to allow for household heterogeneity, which he estimates using information in the third moments. Commault (2022) allows consumption to respond to past transitory shocks and finds a strong response, similar to what experimental studies find. Hryshko and Manovskii (2022) carry out an interesting subsample analysis, identifying sets of households in the PSID with different degrees of insurance based on certain features of their incomes. Crawley and Kuchler (2023) offer another empirical extension, addressing neglected time aggregation in income and consumption in the PSID.
The previous papers restrict income shocks to transmit linearly into consumption. Arellano et al. (2017); Arellano et al. (2023) generalize the income and consumption process by letting permanent shocks feature nonlinear persistence and nonlinear transmission into consumption. [^4] The latter paper advances the former by allowing for flexible heterogeneity and unbalanced panels. Using recent PSID data, they measure the consumption response to income at about 0.3 on average (parameter closest to our definition of pass-through; specification with heterogeneity in the second paper), and the response varies over the distribution of shocks with evidence of asymmetries similar to ours. Identification is quite demanding as they require observing the entire consumption-income joint distribution; our case, by contrast, is far simpler: we need second moments of consumption and up to fourth moments of income. This is appealing in itself (we use more accessible information) and because, as we show later, higher-order features of the consumption distribution exhibit notable discrepancies across survey data, calling into question the conclusions drawn from them. The price of our econometrically more accessible approach is that we have to maintain a linear income process (though one that is subject to tail shocks), while Arellano et al. (2017) do not. Yet, we learn that nonlinearity in income is not crucial for obtaining asymmetric responses to income shocks.
@@
The second branch of the literature studies specific insurance mechanisms through calibrated consumption lifecycle models. [^5] Gourinchas and Parker (2002) and Krueger and Perri (2006) are early precursors in this strand. The early works focus on self-insurance, providing mixed evidence for how much insurance the model permits. Kaplan and Violante (2010) find less insurance compared to BPP, while Guvenen and Smith (2014), who jointly estimate income risk and self-insurance, find more insurance. More recent works study the insurance consequences of labor supply (Heathcote et al., 2014), durable purchases (Madera, 2019), health (Blundell et al., 2020), or added workers (Wu and Krueger, 2021). Except Madera (2019), these papers abstract from higher income moments due to the computational difficulties involved. De Nardi et al. (2019) embed a non-Gaussian income process into a model of consumption/savings and show that tail income risk increases partial insurance due to stronger precautionary motives. In a similar model, Busch and Ludwig (2021) target income skewness and kurtosis, they distinguish between good and bad shocks, and they show that the latter are worse insured. Guvenen et al. (2024) find that tail shocks are worse insured than average shocks, but the benchmark methods of measuring this, effectively BPP, overstate the true extent of insurance. Unlike this quantitative literature, we do not take a stance on the insurance mechanisms available to households; this enables us to measure the overall degree of insurance from all mechanisms together keeping our model partially unspecified, while accounting for higher-order features of the income distribution.
@@
The paper is also related to the newer income dynamics literature that examines the non-Gaussian features of the income distribution. Geweke and Keane (2000) and Bonhomme and Robin (2010) are early works that use the PSID to document, respectively, left skewness and excess kurtosis in earnings growth. Guvenen et al. (2021) use data from the population of working males in the U.S. to provide further evidence of left skewness and excess kurtosis in earnings growth, both accentuated with age and the level of earnings. Guvenen et al. (2014) and Busch et al. (2018) study the cyclical nature of skewness in several countries. While these papers are mostly descriptive of the higher-order features of income, we further document the implications of these features for consumption. We thus contribute to a growing literature on the various economics/finance implications of tail risk, e.g. for growth (Barro, 2009), asset prices (Constantinides and Ghosh, 2017), business cycles (McKay, 2017), consumer welfare (De Nardi et al., 2019), and labor market dynamics (Ai and Bhandari, 2021).
In the remainder of the paper, section 2 presents the model while section 3 establishes identification of the parameters. Section 4 discusses the empirical implementation, section 5 presents the empirical results, and section 6 discusses welfare. Section 7 concludes. The appendix provides theoretical derivations and additional results.
@@ 2 Lifecycle model with tail income risk
where $U(\cdot;\mathbf{Z}_{it})$ is utility over consumption, $\mathbf{Z}_{it}$ is a vector of observed taste shifters (e.g., age, education), $\beta$ is the discount factor, $r$ is the deterministic interest rate, $A_{i0}$ is beginning-of-time financial wealth, ${\operatorname*{\mathbb{E}}}_{0}$ denotes expectations over uncertain future income $Y$, and $T$ is the end of the horizon. We assume that utility satisfies standard regularity conditions, in particular it is continuously twice differentiable, but we do not otherwise parameterize it.

### v4

@@ Abstract
We measure consumption insurance against income shocks accounting for high-order moments of the income distribution. We derive a nonlinear consumption function, where the extent of insurance varies with the shocks’ sign and magnitude. Using PSID data, we estimate an asymmetric pass-through of bad versus good permanent shocks – 17% of a $3\sigma$ negative shock transmits to consumption versus 9% of an equal-sized positive shock – with greater pass-through as shocks worsen. Households would sacrifice over 1/8 of lifetime consumption to eliminate tail income risk. Our results align with survey responses to hypothetical events and suggest that tail risk matters substantially for consumption.
@@ 1 Introduction
In a seminal paper, Blundell et al. (2008) – henceforth BPP – introduce an empirical methodology to measure the pass-through of income shocks of varying persistence into consumption. Using Panel Study of Income Dynamics (PSID) and Consumer Expenditure Survey (CEX) data over 1980–1992, BPP find evidence of partial insurance to permanent shocks and full insurance to transitory ones. BPP rely on a workhorse covariance between income and consumption growth, namely a second moment of their joint distribution. The idea is that the extent to which income growth (driven by income shocks) varies with consumption growth reflects the strength of the transmission of income shocks into consumption.
In a separate strand of literature, Guvenen et al. (2021) use data from the U.S. Social Security Administration and establish important new facts about the distribution of unexplained income growth, namely the distribution of income shocks. They show that income growth is very negatively skewed, i.e., it has a long left tail, so far more people experience large negative than large positive shocks. Moreover, income growth exhibits excess kurtosis, i.e., fat tails, so far more people face either small or large shocks, than moderate ones, relative to a Gaussian density. These are now benchmark facts about income dynamics, observed in many countries (e.g., De Nardi et al., 2021, compare the U.S. and the Netherlands).
If most income shocks away from the mean are large and bad shocks, their effect on consumption might, intuitively, differ from that of small/moderate shocks. In other words, left skewness and excess kurtosis, which are statements about the third and fourth moments of the income distribution respectively, may matter for the pass-through of shocks into consumption. This is precisely what we investigate in this paper. Specifically, we measure the degree of consumption insurance to income shocks of varying persistence, accounting for higher-order moments of their distribution (i.e., tail shocks). This is in contrast to the methodology in BPP, which revolves around the second moments of the joint income-consumption distribution and, thus, abstracts from the modern higher-order facts that characterize income dynamics (the second moment is a poor summary of the tails).
@@
Third, we generalize the method of measuring consumption insurance to explicitly permit a role for tail income risk. BPP log-linearize the policy rule of a lifecycle consumption/savings problem to obtain a consumption function that depends linearly on income shocks. Our first and second contributions above revolve firmly around such a linear framework. We argue (as also does Carroll, 2001) that linearity is a poor approximation to behavior when income is subject to tail shocks. We then derive a higher-order (quadratic) consumption function that allows for nonlinear transmission of shocks. We characterize and empirically measure the consumption pass-through of income shocks, which, in this environment, depends explicitly on the magnitude and sign of shocks. While a nonlinear consumption function has been estimated before, most notably in Arellano et al. (2017) – henceforth ABB, our framework is far less data demanding (we need second moments of consumption and up to fourth moments of income, rather than observing the entire consumption-income joint distribution) and offers a more transparent link with the higher-order features of the distribution of income shocks. Both points have certain important advantages, which we discuss below.
Fourth, we estimate the welfare costs of tail income risk by calculating what fraction of lifetime consumption a household is willing to give up to avoid tail risk. We use parametric utility to calculate these costs, in line with De Nardi et al. (2019), Busch and Ludwig (2024), and Guvenen et al. (2024), but we base these calculations on the pass-through measure from the quadratic consumption function of the previous step. This measures an ‘aggregate’ of all partial insurance available to households, so it is not specific to a particular structural mechanism or a particular parameterization of preferences and expectations.
@@
First, permanent and transitory shocks to disposable household income are negatively skewed, with skewness coefficients at $-0.95$ and $-1.31$, respectively, and highly leptokurtic, with kurtosis coefficients at $44.39$ and $49.23$, respectively. Left skewness implies that a negative shock is typically more unsettling, more hurtful than a positive shock, because it is further away from the zero mean. Excess kurtosis implies that more shocks are concentrated either in the middle of the distribution or far out in the tails, relative to the shoulders. Taken together, large left skewness and excess kurtosis suggest that the realizations of income shocks tend to be either close to their zero mean or far out in the left tail; that is, income shocks for most households are either small/average or quite bad. However, these features vary by age and by the household’s position in the income distribution.
@@
Second, in estimating BPP in the recent years, we find full insurance to transitory shocks and a small pass-through of permanent shocks at 0.15 versus 0.64 in BPP. A 10% permanent income cut thus reduces consumption by only 1.5%, vis-à-vis 6.4% in BPP. In other words, we find much larger partial insurance to permanent shocks in the recent years than originally found in 1980–1992.222Using recent PSID data but a different method, ABB and Arellano et al. (2023) find a similarly low pass-through of permanent shocks. We discuss this further when we present our empirical results. We attribute about a third of the difference from BPP to the consumption imputation from the CEX that BPP use. The imputation brings in large amounts of measurement error, inflating the covariance between income and consumption growth. Another third of the difference arises because the modern PSID provides data biennially (versus annually previously), which tends to mute the transmission of higher frequency shocks. The last third reflects the different calendar times.
@@
Third, introducing the third and fourth moments of income and consumption growth into BPP’s linear consumption setting does not change our estimates of partial insurance. This is counterintuitive in light of the literature that highlights, in multiple settings, the large welfare costs of tail risk (e.g., Barro, 2009; De Nardi et al., 2019). One would thus expect that targeting higher-order moments would reduce the degree of partial insurance, which is not what we find. This result is suggestive of a potential misspecification in the linear consumption function, which motivates our generalization of the methodology through the introduction of a higher-order consumption specification subsequently.
@@
Fourth, in estimating our quadratic consumption function, we find an asymmetry in the pass-through of bad versus good shocks, and of small versus large ones. In this environment, the pass-through depends on the magnitude and sign of shocks; and targeting higher-order income moments enables the identification of the transmission parameters. We draw three main conclusions: 1.) Bad permanent shocks are more unsettling than equal-sized good ones because they have larger transmission rates into consumption – a three standard deviations negative shock (a 50% permanent income cut) has a pass-through of $0.17$ (17% of the shock transmits into consumption), while a positive shock of similar magnitude has a pass-through of $0.09$. 2.) The pass-through of bad shocks increases with the severity of the shock. Bad permanent shocks are thus double hurtful: not only are they more extreme than good ones (they lie further out in the left tail), but they also transmit increasingly more as their magnitude worsens. 3.) Transitory shocks are fully insured in the full sample, at least from a statistical point of view, although the point estimates of their transmission parameters suggest that large transitory windfalls shift consumption considerably, consistent with quasi-experimental evidence of large responses to transitory gains (e.g., Fagereng et al., 2021). Among low wealth low educated households, however, transitory shocks, and in particular negative ones, shift consumption significantly. Overall, the results are consistent with an environment in which households save a good income shock but dissave in the presence of a bad one. There is clearly a limit to how much one can dissave or borrow, so a large negative shock ends up affecting consumption by relatively more.
@@
The asymmetric transmission of bad versus good permanent shocks varies with age and the household’s position in the income distribution. While older households insure average permanent shocks better than younger ones, older households have larger transmission rates of extreme shocks. Severely bad permanent shocks thus cause more harm among the old. Moreover, bad permanent shocks from a given position in the bottom tail of the distribution of income risk cause more harm among the rich (households in the top of the income distribution) than among the poor (households in the bottom); the converse is true for good shocks. Guvenen et al. (2021) show that income skewness becomes more negative with age and income level – there is more room for income to fall among older or higher income workers. In turn, we document that there is also more room for consumption to fall. Among the old, this is because both the rate of transmission and the severity of downwards income risk increase with age; among the rich, however, the asymmetric transmission is driven mostly by the increasing severity of downwards risk along the income distribution.
@@
The seminal paper in the first branch, namely BPP, sparked a large volume of research. [^4] Hall and Mishkin (1982), Deaton and Paxson (1994) and Blundell and Preston (1998) characterize the empirical joint distribution of consumption and income, so they are early precursors in this literature. Blundell et al. (2016) measure partial insurance against wage risk, allowing for family labor supply. Theloudis (2021) allows for preference heterogeneity in partial insurance, which he estimates using information in the third moments. Commault (2022) allows consumption to respond to past transitory shocks and finds a strong response, similar to what experimental studies find. Hryshko and Manovskii (2022) identify sets of households in the PSID with different degrees of insurance based on certain features of their incomes. Crawley and Kuchler (2023) address neglected time aggregation in income and consumption in the PSID.
The previous papers restrict income shocks to transmit linearly into consumption. ABB and Arellano et al. (2023) generalize the income and consumption processes by letting income feature nonlinear persistence, which varies with the level of income and its shock, and nonlinear transmission. Using the recent PSID and a quantile regression approach, they measure the marginal propensity to consume from persistent income, which varies along the income distribution. They simulate the transmission of shocks and document asymmetries between bad/good shocks or between rich/poor, similar to ours. Our papers, however, differ in important dimensions. First, identification in ABB is quite demanding as they require observing the entire consumption-income joint distribution, while we need only second moments of consumption and up to fourth moments of income. This is appealing because we use more accessible information and because, unlike income, we cannot assess the higher-order features of consumption against an established population benchmark. The price of our econometrically more accessible approach is that, unlike ABB, we have to maintain a linear income process (but subject to tail shocks). Yet, we find that nonlinear persistence is not crucial for the main consumption results, i.e., the asymmetric impact of different shocks or at different income levels. Second, ABB measure the pass-through over quantiles of the distribution of income shocks, without explicitly studying the higher-order features of their distribution. We show subsequently that seemingly different quantiles, e.g., shocks in the 10th or 90th percentile as in ABB, may in fact be very similar to one another when the underlying distribution is skewed and leptokurtic. Our model allows us to estimate pass-through rates that explicitly shift with the sign and size of income shocks and can thus describe more transparently how shocks of different signs/sizes shift consumption.
@@
The second branch of the literature studies specific insurance mechanisms through calibrated consumption lifecycle models. [^5] Gourinchas and Parker (2002) and Krueger and Perri (2006) are early precursors in this strand. Early works focus on self-insurance. Kaplan and Violante (2010) find less insurance than BPP, while Guvenen and Smith (2014), who jointly estimate income risk and self-insurance, find more insurance. More recent works study the insurance consequences of labor supply (Heathcote et al., 2014), durables (Madera, 2019), health (Blundell et al., 2020), or added workers (Wu and Krueger, 2021). Except Madera (2019), these papers abstract from higher income moments due to computational difficulties. De Nardi et al. (2019) embed a non-Gaussian income process into a consumption/savings model and show that tail risk increases partial insurance due to stronger precautionary motives. In a similar model, Busch and Ludwig (2024) target income skewness and kurtosis; they distinguish between good and bad shocks and show that the latter are worse insured. Guvenen et al. (2024) find that tail shocks are worse insured than average shocks but the benchmark method in BPP overstates the true extent of insurance. Unlike this quantitative literature, we do not take a stance on the insurance mechanisms available to households; we measure aggregate insurance from all mechanisms together, while accounting for higher-order features of the income distribution, keeping our model parametrically unspecified.
@@
The paper is finally related to the newer income dynamics literature that focuses on non-Gaussian aspects. Geweke and Keane (2000) and Bonhomme and Robin (2010) use the PSID to document, respectively, left skewness and excess kurtosis in earnings growth. Guvenen et al. (2021) use administrative data in the U.S. to provide further evidence of left skewness and excess kurtosis in earnings growth, accentuated with age and the level of earnings. Guvenen et al. (2014) and Busch et al. (2018) study the cyclical nature of skewness in several countries. While these papers are mostly descriptive of the higher-order features of income, we further document the implications of these features for consumption. We thus contribute to a growing literature on the implications of tail risk, e.g., for growth (Barro, 2009), asset prices (Constantinides and Ghosh, 2017), business cycles (McKay, 2017), consumer welfare (De Nardi et al., 2019), and labor market dynamics (Ai and Bhandari, 2021).
In the remainder of the paper, section 2 presents the model and section 3 establishes identification of the parameters. Section 4 discusses the empirical implementation, section 5 presents the results, and section 6 discusses welfare. Section 7 concludes. The supplemental appendix provides additional derivations and results.
@@ 2 Lifecycle model with tail income risk
where $U(\cdot;\mathbf{Z}_{it})$ is utility over consumption, $\mathbf{Z}_{it}$ is a vector of observed taste shifters (e.g., age, education), $\beta$ is the discount factor, $r$ is the deterministic interest rate, $A_{i0}$ is beginning-of-time financial wealth, ${\operatorname*{\mathbb{E}}}_{0}$ denotes expectations over uncertain future income $Y_{it}$, and $T$ is the end of the horizon. We assume that utility satisfies standard regularity conditions, in particular it is continuously twice differentiable, but we do not otherwise parameterize it.